Summary Title: Talking with the Earth
Author’s name: David Mateus
Mission Status: Nominal
Sol Activity Summary: The crew started the day at 8 a.m., to have breakfast and do a communication test with Badalona Comunicacio. At 9 am, the commander and crew engineer received a filming crew from Purdue University. Due to the activities that the filming was doing around the Hab, the crew decided to stay at the science dome most of the time in order to stay in the sim.
The day was focused on the projects and preparing tomorrow’s EVAs. Also, at noon, we had an interview with Badalona Comunicacio.
Look Ahead Plan: Perform two EVA’s
Anomalies in work: None
Weather: Partly cloudy
Crew Physical Status: One of the crew members bled from their nose. You can find more details in the HSO report. The general status of the crew is nominal.
EVA: No EVA today
Reports to be filed: Sol Summary, Journalist report, Photos, Operations report, Green Hab report, Astronomy report, EVA #3 Request, EVA #4 Request, HSO Report.
Support Requested: 9V battery